OpenStudy (anonymous):

show that one and only one out of n,n+2,n+4 is divisible by 3.

OpenStudy (anonymous):

Are you familiar with modular arithmetic?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hm...have you tried anything yet? any ideas on how you might show only one of those numbers can be a multiple of 3?

OpenStudy (anonymous):

ya we have to do this by euclid's algorithm

OpenStudy (anonymous):

Thats an interesting approach. The Euclidean Algorithm helps you find the gcd of 2 numbers. Im not sure how to apply that directly to this problem though =/.

OpenStudy (anonymous):

anyways thks

OpenStudy (anonymous):

One way to do this problem is to note that if you had two multiples of 3 (or multiples of any number for that matter), then when you subtract them, you still have a multiple of 3:\[3m-3n=3(m-n)\]You could use this to get a contradiction.

OpenStudy (anonymous):

If n+2 and n were both multiples of 3, then (n+2)-n should also be a multiple of 3. But...

OpenStudy (anonymous):

if n is even, 2k, 2k+2, 2k+4 2k, 2(k+1), 2(k+2) only one of k,k+1 and k+2 must be divisible by 3. if n is odd, 2k+1, 2k+3, 2k+5 2k+1, 2(k+1)+1, 2(k+2)+1 only one of 2k+1, 2(k+1)+1 and 2(k+2)+1 must be divisible by 3.

ganeshie8 (ganeshie8):

n = 3q + r (euclid's division lemma) n is divisible by 3 only when r = 3k now, for n, n+2, n+4 if we assume one is divisible by 3, others are not divisible by 3 because, they are away by 2 units from r.

OpenStudy (anonymous):

how to write the working?

ganeshie8 (ganeshie8):

here goes : As per Euclid's Division Lemma, Given positive integers a and b, there exist unique integers q and r satisfying a = bq + r, 0 ≤ r < b. a = bq + r since any number of the form 3k is divisible by 3 ( k is an integer) a = 3q + r is divisible by 3 only when r is 0 n, n+2, n+4 are two units away from each other. We can represent any one of these as "a = 3q + r" , (r=0), rest of the two wont satisfy the same because n+2 & n+4 are 2k distance away from each other.

OpenStudy (anonymous):

Simply add the three numbers.. n + n + 2 + n + 4 You will get : 3n + 6 Taking 3 as common, 3(n + 2) this value is divisible by 3 for all values of n because 3 is in multiplication...

OpenStudy (anonymous):

suppose n the number divisible by 3, then we have to show n+2 and n+4 will not be divisible by 3: so n=3p1, where p1 a natural number . by euclid's algorithm, we have: n+2=3x1+2 <since 3 divides n 3=2x1+1 2=1x2 so the gcd of n+2 and 3 is 1, so that 3 cannot divide n+2 n+4=(n+3)+1=3(p1+1)+1< since 3 divides n and ofcourse 3 3=1x3 gcd of n+2 and three is 1 again, so that 3 cannot divide n+2 hence if 3 divides n, it may not divide the others. Now suppose 3 divides n+2 so that n+2=3p2, where p2 is a natural number then: since n+2/3 will give remainder 0, n+1/3 will give remainder 2, then n/3 will give remainder of 1 and then do euclid's algorithm: n=3p3+1 3=1x3 gcd of n and 3 is 1 so that 3 doesn't divide n n+4=(n+3)+1=3(p2+1)+1 3=1x3 so gcd of n+4 and 3 is 1 so that 3 doesn't divide n+4 now we see that if 3 divides n+2, it will not divide the others. And now we suppose 3 divides n+4, so that n+4=3p4, where p4 is a natural number then: since n+4/3 gives remainder 0, n+3/3 gives rem. 2, n+2/3 gives rem. 1, n+1/3 gives rem. 0, n/3 gives rem. 2, so n=3p5+2 3=2x1+1 2=1x2 gcd of n and 3 is 1 so that 3 doesn't divide n n+2=3p6+1 3=1x3 gcd of n+2 and 3 is 1, so that 3 doesn't divide n+2. at last we see that if 3 divides n+4 it will not divide the others. So if 3 divides one of the numbers n,n+2,n+4, it will not divide the others. notice that we could've still proven it without EA, but since you are required to, i've considered it